# Active Sensing for Multiuser Beam Tracking with Reconfigurable Intelligent Surface

### v2

@@ I Introduction
Reconfigurable intelligent surface (RIS) is a promising technology for future wireless communication systems due to its ability to reflect incoming signals toward desired directions in an adaptive fashion [2]. The RIS can enhance communications by establishing focused beams between the access point (AP) and the user equipment (UE), but this focusing capability depends crucially on the availability of channel state information (CSI), which must be obtained in a dedicated channel sensing stage using pilot signals [3, 4, 5]. However, in high-mobility scenarios where CSI needs to be measured frequently, estimating CSI from scratch in each channel sensing phase can lead to large pilot overhead. The main idea of this paper is that significant saving in pilot overhead is possible by exploiting the temporal channel correlation due to UE mobility in the channel sensing stage.
@@
The multiuser beam tracking problem in an RIS-assisted mobile communication system is highly nontrivial, primarily due to the significant pilot overhead required to maintain beam alignment. This is because RIS cannot perform active signal transmission and reception. Consequently, the high-dimensional channels to and from the RIS can only be estimated indirectly based on the low-dimensional observations at the AP. Utilizing active sensing to leverage historical channel information can significantly reduce pilot overhead; but finding the optimal active sensing strategy is very challenging due to the complexity of optimization over a sequence of observations and actions.
@@ I-A Related Works
One of the key differentiating features of this paper as compared to prior works is in the design of active sensing strategies for channel sensing. In most prior works [9, 10, 11, 12, 13, 15, 16, 17], the CSI acquisition stage uses fixed RIS reflection coefficients, either generated randomly or based on fixed patterns such as the discrete Fourier transform (DFT) [25]. However, these non-adaptive RIS sensing schemes are far from optimal. On the other hand, [18] adaptively designs the RIS sensing scheme based on an RNN, but proposes to employ the same RIS reflection coefficients in both sensing and data transmission stages. This is suboptimal due to the different functionalities of the RIS in two stages. A main contribution of this paper is a learning approach for adaptively designing the RIS coefficients in the channel sensing stage for tracking multiple UEs, which to the best of authors’ knowledge has not been explored in the existing literature. Recently, active sensing has demonstrated remarkable performance in related applications including localization [26] and beam alignment [27, 28, 6, 29, 30] in fixed channel environment, where the goal is to leverage previously obtained channel measurements to gradually refine focusing onto some desired low-dimensional part of the static channel. The current investigation is concerned with active sensing for beam tracking in the time-varying channels.
@@ I-B2 Reflection Coefficients Optimization using GNN
This paper proposes to use GNN to learn the mappings from the LSTM cell state vectors to the RIS reflection coefficients for both channel sensing and data transmission. The GNN leverages the spatial relation between the UEs and the RIS to design the reflection patterns while accounting for the mutual interference between the UE’s. The proposed GNN architecture extends the designs in [18, 24, 31], while specifically addressing the distinct roles of the RIS reflection patterns in the sensing vs. data transmission stages. The proposed GNN improves upon previous work [18] that uses the same RIS coefficients for both sensing and communication.
@@ II-A System Model
Consider a narrowband RIS-assisted multiuser multiple-input single-output (MU-MISO) system where $K$ single-antenna user equipment (UEs) are served by an access point (AP) with $M$ antennas. The RIS, equipped with $N_{r}$ passive elements, is placed between the AP and UEs to enable the reflection links as shown in Fig. 1. The RIS works cooperatively with the AP to control the phase of the reflected signals by tuning the phase shifters of each passive element through an RIS controller. We model the time-varying channel as a block-fading channel in discrete time, consisting of a sequence of fixed-length blocks. The channels are constant within the block and are correlated across the successive blocks. We further define a framing structure consisting of $N+1$ blocks per frame, and assume that channel variation is not significant within the frame so a fixed beamforming strategy can be used without incurring substantial performance loss. For the $n$-th block of the $t$-th frame, we use $\bm{G}^{(t,n)}\in\mathbb{C}^{M\times N_{r}}$, $\bm{h}_{\mathrm{d},k}^{(t,n)}\in\mathbb{C}^{M}$, and $\bm{h}_{\mathrm{r},k}^{(t,n)}\in\mathbb{C}^{N_{r}}$ to denote the channel matrix from the RIS to the AP, the channel vector from the UE $k$ to the AP, and the channel vector from the UE $k$ to the RIS, respectively.
@@
In the $n$-th block of the $t$-th frame, let $s_{k}^{(t,n)}\in\mathbb{C}$ be the data symbol transmitted from the AP to the UE $k$, with $\mathbb{E}[|s_{k}^{(t,n)}|^{2}]=1$. The AP transmits $s_{k}^{(t,n)}$ to the UE $k$ through a beamforming vector $\bm{b}_{k}\in\mathbb{C}^{M}$, subject to a power constraint $\sum_{k=1}^{K}\|\bm{b}_{k}^{(t)}\|_{2}^{2}\leq P_{d}$, where $P_{d}$ is the downlink transmit power at the AP. We also denote $\bm{B}^{(t)}=[\bm{b}_{1}^{(t)},\dots,\bm{b}_{K}^{(t)}]\in\mathbb{C}^{M\times K}$ as the beamforming matrix at the AP in the $t$-th frame. Let $\bm{w}^{(t)}=[e^{j\theta_{1,t}},e^{j\theta_{2,t}},\dots,e^{j\theta_{N_{r},t}}]^{\sf{T}}\in\mathbb{C}^{N_{r}}$ be the downlink RIS reflection coefficients in the $t$-th frame, where $\theta_{i,t}\in[0,2\pi)$ is the phase shift of the $i$-th passive RIS element. Then, the signal $y_{k}^{(t,n)}\in\mathbb{C}$ received at the UE $k$ in the $n$-th block of the $t$-th frame can be expressed as:
$y_{k}^{(t,n)}=\sum_{j=1}^{K}(\bm{A}_{k}^{(t,n)}\widetilde{\bm{w}}^{(t)})^{\sf{H}}\bm{b}_{j}^{(t)}s_{j}^{(t,n)}+z_{k}^{(t,n)}\ ,$ (1)
@@
where $\bm{A}_{k}^{(t,n)}\triangleq[\bm{h}_{\mathrm{d},k}^{(t,n)},\bm{G}^{(t,n)}\operatorname{diag}(\bm{h}_{\mathrm{r},k}^{(t,n)})]\in\mathbb{C}^{M\times(N_{r}+1)}$ is the combined channel between the AP and the UE $k$ in the $n$-th block of the $t$-th frame, $\widetilde{\bm{w}}^{(t)}\triangleq[1,(\bm{w}^{(t)})^{\sf{T}}]^{\sf{T}}\in\mathbb{C}^{N_{r}+1}$ is the effective RIS reflection coefficients, and $z_{k}^{(t,n)}\sim\mathcal{CN}(0,\sigma^{2}_{d})$ denotes the i.i.d. AWGN at the UE $k$. Therefore, the downlink achievable data rate $R_{k}^{(t,n)}$ for the UE $k$ in the $n$-th block of the $t$-th transmission frame can be expressed as:
@@
$\mathop{\mathrm{s.t.}}$ ${}_{\mathop{}}\sum_{k=1}^{K}\|\bm{b}_{k}^{(t)}\|_{2}^{2}\leq P_{d}\ ,$
@@ II-B Uplink Pilot Transmission Protocol
Note that there are $(M+N_{r})K+MN_{r}$ unknown channel coefficients in $\bm{G}^{(t,0)}$, $\bm{h}_{r,k}^{(t,0)}$’s, and $\bm{h}_{d,k}^{(t,0)}$’s for $k=1,\dots,K$. Since the number of RIS elements $N_{r}$ is often very large (possibly hundreds), estimating the channels perfectly would lead to significant pilot overhead. It is therefore advantageous to bypass explicit channel estimation and to directly design the beamformers and the RIS reflection patterns based on the received pilots using a neural network [33].
@@ III Active Sensing for Beam Tracking
In (9), we note that the effective channel $\bm{h}_{c,k}^{(t,0)}$ is needed to design the beamforming vectors. Here, since $\bm{h}_{c,k}^{(t,0)}$ is a low-dimensional channel vector, its estimation is straightforward when the RIS reflection coefficients $\widetilde{\bm{w}}^{(t)}$ are fixed. Specifically, this paper simply estimates $\bm{h}_{c,k}^{(t,0)}$ based on the pilot in the $(L+1)$-th sub-block as follows. The decorrelated received pilot $\bar{\bm{y}}_{k,L+1}^{(t)}$ from the UE $k$ in the $(L+1)$-th sub-block is $\bar{\bm{y}}_{k,L+1}^{(t)}=\bm{A}_{k}^{(t,0)}\widetilde{\bm{w}}^{(t)}+\bar{\bm{z}}_{k,L+1}^{(t)}\triangleq\bm{h}_{c,k}^{(t,0)}+\bar{\bm{z}}_{k,L+1}^{(t)}$. So approximately, the effective low-dimensional channel $\check{\bm{h}}_{c,k}^{(t,0)}\in\mathbb{C}^{M}$ is simply:
@@ IV-A GNN for Interference Management
The overall GNN aims to learn the graph representation vector $\bm{s}^{(t)}_{\xi}$’s through an initialization layer, $D$ updating layers, and a final layer that includes normalization. As shown in Fig. 3, the GNN first encodes all the useful information about each node in the $t$-th transmission frame into a corresponding representation vector $\bm{s}^{(t)}_{\xi,0}$’s according to:
@@
To capture the spatial structure between the RIS and UEs, the initial graph representation vector $\bm{s}^{(t)}_{\xi,0}$’s undergo $D$ iterations of updates, each utilizing the output of the preceding layer as input. More specifically, the updates for the RIS node $w$ and $V$ incorporate their current representations as well as the representations of all UE nodes, while the representation vector of the UE node $k$ is updated based on its own prior representation and the representations of all other nodes in the network. These update procedures maintain the GNN’s permutation invariance and equivariance, and therefore allows more efficient training. The updating rules of the $d$-th updating layer are given as:
@@
After $D$ layers of iteration, the updated representation vector $\bm{s}^{(t)}_{\xi,D}$’s would contain sufficient information to design the beam tracking matrices. To obtain the correct output dimensions, we pass $\bm{s}^{(t)}_{k,D}$’s, $\bm{s}^{(t)}_{w,D}$, and $\bm{s}^{(t)}_{V,D}$ through linear layers $g_{1}(\cdot),g_{2}(\cdot),g_{3}(\cdot)$ with $2K$, $2N_{r}$, and $2N_{r}L$ fully connected units, respectively. Specifically, $\bm{s}^{(t)}_{k,D+1}=g_{1}(\bm{s}^{(t)}_{k,D})\in\mathbb{R}^{2K}$, $\bm{s}^{(t)}_{w,D+1}=g_{2}(\bm{s}^{(t)}_{w,D})\in\mathbb{R}^{2N_{r}}$, and $\bm{s}^{(t)}_{V,D+1}=g_{3}(\bm{s}^{(t)}_{V,D})\in\mathbb{R}^{2N_{r}L}$. The final representation vector $\bm{s}^{(t)}_{\xi,D+1}$’s are then normalized to produce the RIS sensing vectors $\{\bm{v}^{(t+1)}_{\ell}\}^{L}_{\ell=1}$, downlink RIS reflection coefficients $\bm{w}^{(t)}$, and power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$, respectively, while ensuring that the unit modulus constraints on $\bm{w}^{(t)}$, $\bm{v}_{\ell}^{(t+1)}$’s and the power constraints on $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$ are satisfied.
@@ IV-B LSTM for Exploiting Temporal Channel Correlation
This paper proposes to use an LSTM-based RNN to capture the temporal correlations of the time-varying channel by automatically summarizing the channel information into state information vector $\bm{c}^{(t)}_{k}$’s in (13), which are further processed by the GNN to design the beam tracking strategies. As shown in Fig. 4, $K$ copies of the LTSM cells with shared weights are deployed to summarize the state information for the $K$ mobile UEs, respectively. In the $t$-th frame, the $k$-th LSTM cell takes new pilots $\bar{\bm{Y}}_{k}^{(t)}$ from UE $k$ to update its hidden state vector $\bm{h}_{k}^{(t)}$ and cell state vector $\bm{c}_{k}^{(t)}$ according to [7]:
@@
Figure 5: Proposed active sensing framework for multiuser beam tracking with RIS. During the training phase, we concatenated $U$ active sensing units correspond to $U$ transmission frames. Once trained and deployed, we reuse the active sensing units for a potentially infinite number of transmission frames.

### v3

@@ I Introduction
Reconfigurable intelligent surface (RIS) is a promising technology for future wireless communication systems due to its ability to reflect incoming signals toward desired directions in an adaptive fashion [2]. The RIS can enhance communications by establishing focused beams between the access point (AP) and the user equipment (UE), but this focusing capability depends crucially on the availability of channel state information (CSI), which must be obtained in a dedicated channel sensing stage using pilot signals [3, 4, 5]. However, in high-mobility scenarios where CSI needs to be measured frequently, estimating CSI from scratch in each channel sensing phase can lead to a large pilot overhead. The main idea of this paper is that significant saving in pilot overhead is possible by exploiting the temporal channel correlation due to UE mobility in the channel sensing stage.
@@
The multiuser beam tracking problem in an RIS-assisted mobile communication system is highly nontrivial, primarily due to the significant pilot overhead required to maintain beam alignment. This is because RIS cannot perform active signal transmission or reception. Consequently, the high-dimensional channels to and from the RIS can only be estimated indirectly based on the low-dimensional observations at the AP. Utilizing active sensing to leverage historical channel information can significantly reduce pilot overhead; but finding the optimal active sensing strategy is very challenging due to the complexity of optimization over a sequence of observations and actions.
@@ I-A Related Works
One of the key differentiating features of this paper as compared to prior works is in the design of active sensing strategies for channel sensing. By active sensing, we mean that the AP adaptively and sequentially queries the environment to measure and to track the channel dynamics, by using a sequence of focused sensing beams towards the mobile UEs. In particular, at each CSI acquisition stage, the AP adaptively designs the sensing patterns at the RIS based on the information obtained from all the previous stages. In most prior works [9, 10, 11, 12, 13, 15, 16, 17], the CSI acquisition stage uses fixed RIS reflection coefficients, either generated randomly or based on fixed patterns such as the discrete Fourier transform (DFT) [25]. However, these non-adaptive RIS sensing schemes are far from optimal. On the other hand, [18] adaptively designs the RIS sensing scheme based on an RNN, but proposes to employ the same RIS reflection coefficients in both sensing and data transmission stages. This is suboptimal due to the different functionalities of the RIS in two stages. A main contribution of this paper is a learning approach for adaptively designing the RIS coefficients in the channel sensing stage for tracking multiple UEs, which to the best of authors’ knowledge has not been explored in the existing literature. Recently, active sensing has demonstrated remarkable performance in related applications including localization [26] and beam alignment [27, 28, 6, 29, 30] in fixed channel environment, where the goal is to leverage previously obtained channel measurements to gradually refine focusing onto some desired low-dimensional part of the static channel. The current investigation is concerned with active sensing for beam tracking in the time-varying channels.
@@ I-B2 Reflection Coefficients Optimization Using GNN
This paper proposes to use GNN to learn the mappings from the LSTM cell state vectors to the RIS reflection coefficients for both channel sensing and data transmission. The GNN leverages the spatial relation between the UEs and the RIS to design the reflection patterns while accounting for the mutual interference between the UEs. The proposed GNN architecture extends the designs in [18, 24, 31], while specifically addressing the distinct roles of the RIS reflection patterns in the sensing vs. data transmission stages. The proposed GNN improves upon previous work [18] that uses the same RIS coefficients for both sensing and communication.
@@ II-A System Model
Consider a narrowband RIS-assisted multiuser multiple-input single-output (MU-MISO) system where $K$ single-antenna user equipment (UEs) are served by an access point (AP) with $M$ antennas. The RIS, equipped with $N_{r}$ passive elements, is placed between the AP and UEs to enable the reflection links as shown in Fig. 1. The RIS is assumed to operate under ideal conditions with perfect phase control and negligible hardware impairments. The RIS controller works cooperatively with the AP to control the magnitudes and directions of the reflected signals by tuning the phase shifters of the passive elements of the RIS. We model the time-varying channel in the discrete-time domain, where the continuous evolution of the channel is approximated by dividing time into a sequence of fixed-length blocks. The channel coefficients are constant within the block and are correlated across the successive blocks. This correlation reflects the gradual changes in the channel due to the mobility of the UEs [32, 33]. We further define a framing structure consisting of $N+1$ blocks per frame, and assume that channel variation is not significant within the frame so a fixed beamforming strategy across all the blocks within the frame can be used without incurring substantial performance loss. For the $n$-th block of the $t$-th frame, we use $\bm{G}^{(t,n)}\in\mathbb{C}^{M\times N_{r}}$, $\bm{h}_{\mathrm{d},k}^{(t,n)}\in\mathbb{C}^{M}$, and $\bm{h}_{\mathrm{r},k}^{(t,n)}\in\mathbb{C}^{N_{r}}$ to denote the channel matrix from the RIS to the AP, the channel vector from the UE $k$ to the AP, and the channel vector from the UE $k$ to the RIS, respectively.
@@
In the $n$-th block of the $t$-th frame, let $s_{k,i}^{(t,n)}\in\mathbb{C}$ be the $i$-th data symbol transmitted from the AP to the UE $k$, with $\mathbb{E}[|s_{k,i}^{(t,n)}|^{2}]=1$. The AP transmits $s_{k,i}^{(t,n)}$ to the UE $k$ through a beamforming vector $\bm{b}_{k}^{(t)}\in\mathbb{C}^{M}$, subject to a power constraint $\sum_{k=1}^{K}\|\bm{b}_{k}^{(t)}\|_{2}^{2}\leq P_{d}$, where $P_{d}$ is the downlink transmit power at the AP. We use $\bm{B}^{(t)}=[\bm{b}_{1}^{(t)},\dots,\bm{b}_{K}^{(t)}]\in\mathbb{C}^{M\times K}$ to denote the beamforming matrix at the AP in the $t$-th frame. Let $\bm{w}^{(t)}=[e^{j\theta_{1,t}},e^{j\theta_{2,t}},\dots,e^{j\theta_{N_{r},t}}]^{\sf{T}}\in\mathbb{C}^{N_{r}}$ be the downlink RIS reflection coefficients in the $t$-th frame, where $\theta_{u,t}\in[0,2\pi)$ is the phase shift of the $u$-th passive RIS element. Then, the signal $y_{k,i}^{(t,n)}\in\mathbb{C}$ received at the UE $k$ in the $n$-th block of the $t$-th frame can be expressed as:
$y_{k,i}^{(t,n)}=\sum_{j=1}^{K}(\bm{A}_{k}^{(t,n)}\widetilde{\bm{w}}^{(t)})^{\sf{H}}\bm{b}_{j}^{(t)}s_{j,i}^{(t,n)}+z_{k,i}^{(t,n)},$ (1)
@@
where $\bm{A}_{k}^{(t,n)}\triangleq[\bm{h}_{\mathrm{d},k}^{(t,n)},\bm{G}^{(t,n)}\operatorname{diag}(\bm{h}_{\mathrm{r},k}^{(t,n)})]\in\mathbb{C}^{M\times(N_{r}+1)}$ is the combined channel between the AP and the UE $k$ in the $n$-th block of the $t$-th frame, $\widetilde{\bm{w}}^{(t)}\triangleq[1,(\bm{w}^{(t)})^{\sf{T}}]^{\sf{T}}\in\mathbb{C}^{N_{r}+1}$ is the effective RIS reflection coefficients, and $z_{k,i}^{(t,n)}\sim\mathcal{CN}(0,\sigma^{2}_{d})$ denotes the i.i.d. Gaussian noise at the UE $k$. Therefore, the downlink achievable data rate $R_{k}^{(t,n)}$ for the UE $k$ in the $n$-th block of the $t$-th transmission frame can be expressed as:
@@
$\mathop{\mathrm{subject\ to}}$ ${}_{\mathop{}}\sum_{k=1}^{K}\|\bm{b}_{k}^{(t)}\|_{2}^{2}\leq P_{d}$
@@ II-B Uplink Pilot Transmission Protocol
Note that there are $(M+N_{r})K+MN_{r}$ unknown channel coefficients in $\bm{G}^{(t,0)}$, $\bm{h}_{r,k}^{(t,0)}$’s, and $\bm{h}_{d,k}^{(t,0)}$’s for $k=1,\dots,K$. Since the number of RIS elements $N_{r}$ is often very large (possibly hundreds), estimating the channels perfectly would lead to a significant pilot overhead. It is therefore advantageous to bypass explicit channel estimation and to directly design the beamformers and the RIS reflection patterns based on the received pilots using a neural network [35].
@@ III Active Sensing for Beam Tracking
In (9), we note that the effective channel $\bm{h}_{c,k}^{(t,0)}$ is needed to design the beamforming vectors. Here, since $\bm{h}_{c,k}^{(t,0)}$ is a low-dimensional channel vector, its estimation is straightforward when the RIS reflection coefficients $\widetilde{\bm{w}}^{(t)}$ are fixed. Specifically, this paper simply estimates $\bm{h}_{c,k}^{(t,0)}$ based on the pilot in the $(L+1)$-th sub-block as follows. The decorrelated received pilot $\bar{\bm{y}}_{k,L+1}^{(t)}$ from the UE $k$ in the $(L+1)$-th sub-block is $\bar{\bm{y}}_{k,L+1}^{(t)}=\bm{A}_{k}^{(t,0)}\widetilde{\bm{w}}^{(t)}+\bar{\bm{z}}_{k,L+1}^{(t)}\triangleq\bm{h}_{c,k}^{(t,0)}+\bar{\bm{z}}_{k,L+1}^{(t)}$. So approximately, the effective low-dimensional channel $\check{\bm{h}}_{c,k}^{(t,0)}\in\mathbb{C}^{M}$ is simply
@@ IV-A GNN for Interference Management
The overall GNN aims to learn the graph representation vectors $\bm{s}^{(t)}_{\xi}$ through an initialization layer, $D$ updating layers, and a final layer that includes normalization. As shown in Fig. 3, the GNN first encodes all the useful information about each node in the $t$-th transmission frame into a corresponding representation vector $\bm{s}^{(t)}_{\xi,0}$ according to:
@@
To capture the spatial structure between the RIS and UEs, the initial graph representation vectors $\bm{s}^{(t)}_{\xi,0}$ undergo $D$ iterations of updates, each utilizing the output of the preceding layer as input. More specifically, the updates for the RIS node $w$ and $V$ incorporate their current representations as well as the representations of all UE nodes, while the representation vector of the UE node $k$ is updated based on its own prior representation and the representations of all other nodes in the network. These update procedures maintain the GNN’s permutation invariance and equivariance, and therefore allows more efficient training. The updating rules of the $d$-th updating layer are given as:
@@
After $D$ layers of iteration, the updated representation vectors $\bm{s}^{(t)}_{\xi,D}$ would contain sufficient information to design the beam tracking matrices. To obtain the correct output dimensions, we pass $\bm{s}^{(t)}_{k,D}$’s, $\bm{s}^{(t)}_{w,D}$, and $\bm{s}^{(t)}_{V,D}$ through linear layers $g_{1}(\cdot),g_{2}(\cdot),g_{3}(\cdot)$ with $2$, $2N_{r}$, and $2N_{r}L$ fully connected units, respectively. Specifically, $\bm{s}^{(t)}_{k,D+1}=g_{1}(\bm{s}^{(t)}_{k,D})\in\mathbb{R}^{2}$, $\bm{s}^{(t)}_{w,D+1}=g_{2}(\bm{s}^{(t)}_{w,D})\in\mathbb{R}^{2N_{r}}$, and $\bm{s}^{(t)}_{V,D+1}=g_{3}(\bm{s}^{(t)}_{V,D})\in\mathbb{R}^{2N_{r}L}$. The final representation vectors $\bm{s}^{(t)}_{\xi,D+1}$ are then normalized to produce the RIS sensing vectors $\{\bm{v}^{(t+1)}_{\ell}\}^{L}_{\ell=1}$, downlink RIS reflection coefficients $\bm{w}^{(t)}$, and power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$, respectively, while ensuring that the unit modulus constraints on $\bm{w}^{(t)}$, $\bm{v}_{\ell}^{(t+1)}$’s and the power constraints on $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$ are satisfied.
@@ IV-B LSTM for Exploiting Temporal Channel Correlation
This paper proposes to use an LSTM-based RNN to capture the temporal correlations of the time-varying channel by automatically summarizing the channel information into state information vectors $\bm{c}^{(t)}_{k}$ in (13), which are further processed by the GNN to design the beam tracking strategies. As shown in Fig. 4, $K$ copies of the LTSM cells with shared weights are deployed to summarize the state information for the $K$ mobile UEs, respectively. In the $t$-th frame, the $k$-th LSTM cell takes new pilots $\bar{\bm{Y}}_{k}^{(t)}$ from UE $k$ to update its hidden state vector $\bm{h}_{k}^{(t)}$ and cell state vector $\bm{c}_{k}^{(t)}$ according to [7]:
@@
Figure 5: Proposed active sensing framework for multiuser beam tracking with RIS. During the training phase, we concatenate $U$ active sensing units corresponding to $U$ transmission frames. Once trained and deployed, we reuse the active sensing units for a potentially infinite number of transmission frames.